# Tracing the Techno-Supremacy Doctrine: A Critical Discourse Analysis of the AI Executive Elite

### v1

@@ 5.2 The Elite’s Discursive Shift
The social context discussed in the previous section has influenced (and been influenced by) the discourse of the AI executive elite. This section examines potential social determinants contributing to the recent surge in pro-TSD discourse. This analysis is framed in terms of the technological, commercial, and geopolitical forces outlined in the previous section.
@@
First, the increasingly impressive and frequent technological breakthroughs of the last few years likely fueled a sense of awe (CT-MP) and inspired visions of shared prosperity (CT-UF). Four of the top five CT-MP texts and four of the top five CT-UF texts were published within this period. Second, the AI gold rush, fueled by immense commercial interests, likely led to perceiving AI as the next engine of economic growth (TC-PD). Technology companies have been rapidly adding AI capabilities to virtually all their products (cf. Tech Goggles, Solutionism), claiming significant gains in productivity (TG-PE). Four of the top five Tech Goggles texts and the top five Solutionism texts were published within this period. Finally, the current hyper-competition has been fueled by the US administration’s vocal support for American leadership as a reaction to China’s AI development. The “bad actor” threat is commonly used to justify many AI-related concerns (ADD-JU). Three of the top five ADD-JU texts were published within this period.
@@
The determinants discussed thus far have likely contributed to a surge in the expression of core TSD tenets (cf. TCE codes). However, the period following the launch of ChatGPT also saw a noticeable surge in TSD-reinforcing responses to AI-related concerns (cf. TRR codes), particularly those devoted to marginalization. This surge in TRR discourse can be primarily explained by discursive, rather than technological, commercial, or geopolitical determinants. Ironically, the dominance of the safety-first discourse around the time ChatGPT was launched may have led to the marginalization of concerns. Most elite members (18 of 19) indicated their divergence from the dominant safety discourse by not signing the Future of Life Institute’s open letter in early 2023. Shortly after, this implicit stance became explicit opposition as many members started minimizing, dismissing, deflecting, or reframing AI-related concerns (cf. MAR codes). Indeed, four of the top five MAR texts were published within this period. The most extreme example of this overt conflict is andreessen-2023-manifesto, a manifesto with an explicit call to action against a perceived “enemy,” displaying the most voluminous dismissive discourse (MAR-DI) across the corpus. These observations suggest that discourse is not merely a result of its evolving social context, but a part of it, meaning that it can also be a powerful determinant for future discourse.
@@ 6.1 Key Findings
In response to objective O1, this research demonstrates conclusively that the TSD discourse of the AI executive elite is not monolithic but reveals a broad spectrum of stances. While the average text tends to be moderately pro-TSD, this central tendency is marked by high variability, confirming that TSD is not a uniformly held belief system but a site of active discursive struggle. The analysis reveals that pro-TSD discourse is predominantly constructed through direct expressions of its core tenets, particularly visions of a utopian future and the framing of technology as the primary or inevitable driver of progress. Moreover, these two aspects of TSD are often combined, creating a powerful legitimizing narrative for the accelerated development and deployment of AI at scale. A key finding is the identification of Benign Techno-Optimism (BTO) and its distinct discursive profile. While displaying a genuine belief in AI’s positive potential, BTO discourse is critically tempered by direct contradictions to core TSD tenets, particularly those related to Techno-Chauvinism, and explicitly values non-technological efforts. The distinction between full-fledged TSD and BTO ultimately revolves around their implicit theories of change. TSD posits that technology is the primary, autonomous driver of progress, with society needing to adapt accordingly. In contrast, BTO frames technology as a powerful but contingent tool that must be actively and deliberately steered by humanistic values, governance, and social institutions to produce beneficial outcomes. The struggle between these two visions is not merely about optimism versus pessimism, but about where agency and ultimate control reside, in the technology itself or in the society that wields it. BTO’s optimism is earned, not assumed; it coexists with a deep awareness of risks and an openness to diverse, non-technological mitigation strategies. While BTO may be less problematic than full-fledged TSD, it is crucial to critically question whether it could function as a more sophisticated form of legitimation that preserves the industry’s central role under the more palatable guise of ‘responsible innovation.’ BTO discourse could serve as a strategic shield, deflecting more radical critiques and ensuring the industry remains the primary architect of its own oversight.
@@
Regarding objective O2, the research finds that the most common form of critical engagement is not direct opposition to TSD core tenets but the acknowledgment of AI-related concerns. However, a crucial finding is that this acknowledgment frequently functions as a mere rhetorical preamble. It strategically validates concerns only to pivot to pro-TSD responses, most notably by proposing purely technological solutions or marginalizing the concern. This pattern reveals that while elite members frequently engage with risk, they rarely pose fundamental challenges to TSD itself.
@@
In response to objective O3, the analysis reveals that the discourse of the AI executive elite is not static but highly dynamic and adaptive. A key finding is the marked polarization of stances after the launch of ChatGPT, with most authors exhibiting a notable increase in pro-TSD discourse over time. This discursive shift seems to be a strategic adaptation to a rapidly evolving social context, moving from an initial focus on AI safety to action. This dynamic can be explained by a confluence of social determinants: impressive technological breakthroughs fueled the tenets of Techno-Chauvinism and the Cult of Technology; intense commercial competition promoted Solutionist and Tech Goggles framings; and geopolitical pressures provided justification for accelerated development. The increase in discourse marginalizing AI-related concerns seems to be a direct counter-narrative to the rising influence of the previously dominant safety-first discourse. This phenomenon positions the AI executive elite’s not merely as a reaction to the ‘AI gold rush,’ but as a strategic intervention that actively works to construct, legitimize, and accelerate concrete visions for the future of AI.
@@ 6.2 A Diagnostic Approach
TSD, with its unfortunate anagram, can be thought of as an ideological contagion that spreads through discourse. Like a pathogen, its influence is not always immediately apparent. It can be challenging to detect because its tenets are often naturalized through repetition, eventually appearing as common sense. TSD can thus ‘infect’ even the most well-intentioned efforts to tackle important issues, subtly shaping the framing of problems and limiting the range of considered solutions.
@@
Given this often subtle nature and the broad range of manifestations, developing a suitable diagnostic toolkit for identifying TSD becomes crucial. The methodology developed for this study functions like such a toolkit. It aims to move beyond intuition, offering a systematic method to detect TSD’s presence, much like a trained medical professional uses specific tests to confirm a diagnosis rather than relying on general impressions. The quantitative metrics (TSD Adherence and Balance) provide quick tests, offering a rapid overview of a text to uncover broad patterns, identify outliers, and articulate well-supported hypotheses. These quantitative insights inform the in-depth qualitative analyses that serve as more specialized diagnostic tests. These in-depth analyses examine the specific linguistic ‘symptoms’ to confirm the presence of TSD and understand its distinct manifestation (e.g., benign, insidious, aggressive). Just as an experienced doctor becomes progressively better at spotting diseases, TSD becomes easier to trace and critique once an analyst is sensitized to its discursive markers.
The use of this metaphor is conscious and intentional. Metaphors are not merely literary and discursive tools; they are used in people’s thought processes and, consequently, influence their actions (Lakoff and Johnson, 2003). Beyond illustrative purposes, the TSD diagnostic toolkit metaphor is intended to be drawn upon frequently when producing and interpreting discourse related to technology and its impact on society.
6.3 From Diagnosis to Praxis
For CDA to contribute to social change, it must engage in dialogue with social actors who are in a position to act. The ultimate value of this study lies in fostering awareness among key stakeholders. This awareness, however, must be critical and extend beyond simply being able to trace TSD. It is also crucial for these social actors to understand why it can be problematic and how its components reinforce one another. There are three social groups that must develop this critical awareness for TSD to be challenged effectively: AI practitioners, policymakers, and consumers.
@@ 6.3 From Diagnosis to Praxis
AI practitioners building and deploying applications (e.g., engineers, designers) must be able to diagnose TSD in their discourse and that of their leadership. Without this critical awareness, even well-intentioned efforts can fall into the techno-solutionist trap. By becoming sensitized to TSD, AI practitioners can question whether a technological fix is the best approach and implement genuine solutions that account for complex social contexts and integrate appropriate safeguards. Policymakers and regulators formulating governance frameworks (e.g., the EU AI Act) are subject to intense lobbying and sophisticated public discourse. Being able to detect the subtle workings of TSD (e.g., distinguishing a genuine acknowledgment of risk from a token gesture meant to preempt regulation) is essential for crafting effective, robust, and truly society-serving policies that are not captured by the very industry they seek to govern. AI consumers (i.e., virtually everyone with access to digital technology) have a crucial and often underestimated role. If consumers are sensitized to TSD, they can purposefully exercise their consumer and civic power to question assumptions and claims that often dominate marketing and public narratives.
@@
Developing a critical awareness to trace TSD is necessary but insufficient for meaningful praxis. While the diagnostic toolkit developed in this study empowers social actors to identify and support more responsible manifestations (e.g., Benign Techno-Optimism) over more insidious or aggressive ones, the final stage of critical engagement requires holding elite members accountable by verifying the congruence between their public discourse and their material actions. Accountability demands that we scrutinize whether a company’s stated commitment to safety (e.g., Anthropic’s Responsible Scaling Policy) is reflected in its product design choices; whether its advocacy for responsible AI aligns with its lobbying efforts on regulation; and whether its vision of shared prosperity matches its internal labor practices and investment priorities. Without this vital step of verifying alignment between word and deed, even the most seemingly responsible discourse risks functioning merely as a strategic shield for unaccountable power.
@@ 7 Conclusion
7.1 Limitations and Future Work
The contributions above should be contextualized by the methodological limitations already discussed and additional broader limitations discussed in this section. These boundaries serve as potential avenues for future research.
First, TSD may not be a consciously held belief system. While this study defines TSD as a cohesive ideology, its tenets might be more fragmented or opportunistic in practice. Future ethnographic or interview-based research could explore the extent to which this framework aligns with the self-perceptions and motivations of elite members. Moreover, it could investigate the out-of-scope phenomena identified during the broad phase of this study, including the role of saviorism given perceived moral motivations; internal disagreements and fragmentation within the elite (e.g., Andreessen versus Amodei on the merits of Universal Basic Income); and the quasi-religious nature of TSD discourse.
A second limitation is the study’s scope. The focus on a systematically defined, US-centric executive elite means the findings may not generalize to the broader global AI ecosystem. Future analyses could include a broader range of influential actors, including leaders of non-US companies (e.g., DeepSeek, Mistral) and key opinion formers (e.g., scientists, journalists). These analyses would create a comparative understanding of how TSD is reproduced or challenged globally and across different institutional roles. Additionally, the scope could be broadened to include shorter texts (e.g., tweets) and spoken discourse (e.g., podcasts), allowing for the inclusion of highly influential figures like Demis Hassabis and Elon Musk.
Another limitation is that this study does not establish a causal relationship between elite discourse and societal outcomes. While premised on the principle that discourse has social effects, it cannot isolate or measure the actual impact of these discursive strategies in the real world. This limitation points to the need for empirical reception studies. Future projects could employ experimental methods or surveys to investigate how different manifestations of TSD, from benign to insidious, actually influence public opinion or the attitudes of policymakers.
@@ 7.1 Limitations and Future Work
The most personally exciting and immediate future project is the development of an AI-powered TSD diagnostic tool. The underlying model would be fine-tuned on the manual annotations and findings, and prompted to act as an expert in this study’s methodology. The model could employ Retrieval-Augmented Generation (RAG) techniques to have access to TSD’s and CDA’s theoretical foundations and utilize Reinforcement Learning (RL) to incorporate user feedback on an ongoing basis. Such an application could analyze text and audio in near real-time, providing users with TSDA and TSDB scores, and identify concrete linguistic features that signal TSD’s presence. To maximize its practical impact, the tool could position analyzed texts on the TSDA/TSDB spectrum relative to this study’s corpus and, crucially, link identified TSD aspects to established critiques and concrete examples of potential or actual harm. By presenting these insights in accessible formats (e.g., infographics, short videos, podcasts), this tool could translate critical academic research into a powerful instrument for public and professional awareness. This project would represent a reflexive application of AI, turning its analytical power back upon the ideological systems from which it emerges. This approach moves beyond a simple anti-technology stance to embody a form of critical praxis. By leveraging the unique capabilities of AI for scalable, real-time analysis, the goal is not to propose a technological ‘fix’ for TSD, but to create an instrument for public sensitization.
7.2 Take-Home Message
@@
This study is a piece of anti-TSD discourse from a concerned techno-optimist. While technology can be a powerful aid to tackling complex societal issues, it must be employed consciously to equitably prevent or mitigate harm. The sheer velocity of AI development with its accompanying legitimizing discourse makes critical awareness an immediate necessity. While awareness alone is not transformative action, it is an essential prerequisite. By uncovering the Techno-Supremacy Doctrine’s core components, discussing its risks, and providing a diagnostic toolkit for its manifestations, this study aims to foster a collective sensitization among AI practitioners, policymakers, and consumers.
We all have the responsibility to act. Let us support or challenge dominant discourses both consciously and materially, demanding that the promises made are met with accountable action. Let us navigate the future of AI as active participants, not passive recipients of an “inevitable” destiny manufactured unilaterally by the AI executive elite.

### v2

@@ 4.2.2 The Elite’s Discursive Shift
The social context discussed in the previous section has influenced (and been influenced by) the discourse of the AI executive elite. This section examines potential social determinants contributing to the recent surge in pro-TSD discourse. This analysis is framed by the technological, commercial, and geopolitical forces outlined in the previous section.
@@
First, the increasingly impressive and frequent technological breakthroughs of the last few years likely fueled a sense of awe and inspired visions of shared prosperity. Four of the top five (80%) texts displaying Cult of Technology aspects were published within this period. Second, the AI ‘gold rush,’ fueled by immense commercial interests, likely led to viewing AI as the next engine of economic growth. Technology companies have been rapidly adding AI capabilities to virtually all their products (cf. Tech Goggles, Solutionism), claiming significant gains in productivity. Four of the top five (80%) Tech Goggles texts and the top five Solutionism texts were published within this period. Finally, the current hyper-competition has been fueled by the US administration’s vocal support for American leadership in response to China’s AI development. The “bad actor” threat is commonly used to justify many AI-related concerns. Three of the top five (60%) texts were published within this period.
@@
The determinants discussed thus far have likely contributed to a surge in the expression of core TSD tenets. However, the period following the launch of ChatGPT also saw a noticeable surge in TSD-reinforcing responses to AI-related concerns, particularly those devoted to marginalization. This surge can be primarily explained by discursive, rather than technological, commercial, or geopolitical determinants. Ironically, the dominance of the safety-first discourse around the time ChatGPT was launched may have led to the marginalization of concerns. In early 2023, 18 out of 19 (95%) elite members indicated their divergence from the dominant safety discourse by not signing the Future of Life Institute’s open letter. Shortly after, this implicit stance became explicit opposition as many members started minimizing, dismissing, deflecting, or reframing AI-related concerns. Indeed, four of the top five (80%) marginalizing texts were published within this period. The most extreme example of this overt conflict is andreessen-2023-manifesto, a manifesto with an explicit call to action against a perceived “enemy,” and the most voluminous source of dismissive discourse across the corpus. These observations suggest that discourse is not merely a result of its evolving social context, but a part of it, meaning that it can also be a powerful determinant for future discourse.
@@ 5 Discussion
In response to objective I (how TSD is presented, challenged, or reproduced), this research demonstrates conclusively that the TSD discourse of the AI executive elite is not monolithic but reveals a broad spectrum of stances. While the average text tends to be moderately pro-TSD, this central tendency is marked by high variability, confirming that TSD is not a uniformly held belief system but a site of active discursive struggle. The analysis reveals that pro-TSD discourse is predominantly constructed through direct expressions of its core tenets, particularly visions of a utopian future and the framing of technology as the primary or inevitable driver of progress. Moreover, these two aspects of TSD are often combined, creating a powerful legitimizing narrative for the accelerated development and deployment of AI at scale. A key finding is the identification of Benign Techno-Optimism (BTO) and its distinct discursive profile. While displaying a genuine belief in AI’s positive potential, BTO discourse is critically tempered by direct contradictions to core TSD tenets, particularly those related to Techno-Chauvinism, and explicitly values non-technological efforts. The distinction between full-fledged TSD and BTO ultimately revolves around their implicit theories of change. TSD posits that technology is the primary, autonomous driver of progress, and that society must adapt accordingly. In contrast, BTO frames technology as a powerful but contingent tool that must be actively and deliberately steered by humanistic values, governance, and social institutions to produce beneficial outcomes. The struggle between these two visions is not merely about optimism versus pessimism, but about where agency and ultimate control reside, in the technology itself or in the society that wields it. BTO’s optimism is earned, not assumed; it coexists with a deep awareness of risks and an openness to diverse, non-technological mitigation strategies. While BTO may be less problematic than full-fledged TSD, it is crucial to critically question whether it could function as a more sophisticated form of legitimation that preserves the industry’s central role under the more palatable guise of ‘responsible innovation.’ BTO discourse could serve as a strategic shield, deflecting more radical critiques and ensuring the industry remains the primary architect of its own oversight.
@@
Regarding objective II (response to AI-related concerns), the research finds that the most common form of critical engagement is not direct opposition to TSD core tenets but the acknowledgment of concerns. However, a crucial finding is that this acknowledgment often serves merely as a rhetorical preamble. It strategically validates concerns only to pivot to pro-TSD responses, most notably by proposing purely technological solutions or marginalizing the concern. This pattern reveals that while elite members frequently engage with risk, they rarely pose fundamental challenges to TSD itself.
@@
In response to objective III (elite’s discourse dynamics), the analysis reveals that the AI executive elite’s discourse is not static but highly dynamic and adaptive. A key finding is the marked polarization of stances after the launch of ChatGPT, with most authors exhibiting a notable increase in pro-TSD discourse over time. This discursive shift seems to be a strategic adaptation to a rapidly evolving social context, moving from an initial focus on AI safety to action. This dynamic can be explained by a confluence of social determinants: impressive technological breakthroughs fueled the tenets of Techno-Chauvinism and the Cult of Technology; intense commercial competition promoted Solutionist and Tech Goggles framings; and geopolitical pressures provided justification for accelerated development. The increase in discourse marginalizing AI-related concerns seems to be a direct counter-narrative to the rising influence of the previously dominant safety-first discourse. This phenomenon positions the AI executive elite’s discourse not merely as a reaction to the ‘AI gold rush,’ but as a strategic intervention that actively works to construct, legitimize, and accelerate concrete visions for the future of AI.
@@ 5.1 A Diagnostic Approach
TSD (with its unfortunate anagram) can be thought of as an ideological contagion that spreads through discourse. Like a pathogen, its influence is not always immediately apparent. It can be challenging to detect because its tenets are often naturalized through repetition, eventually appearing as common sense. TSD can thus ‘infect’ even the most well-intentioned efforts to tackle important issues, subtly shaping the framing of problems and limiting the range of considered solutions. Given its often subtle nature and the broad range of manifestations, developing a suitable diagnostic toolkit to identify TSD becomes crucial. The methodology developed for this paper functions like such a toolkit. It aims to move beyond intuition, offering a systematic methodology for detecting TSD, much as a trained medical professional uses specific tests to confirm a diagnosis rather than relying on general impressions. The quantitative metrics (TSD Adherence and Balance) provide quick tests that offer a rapid overview of a text, uncovering broad patterns and identifying outliers to articulate well-supported hypotheses. These quantitative insights inform the in-depth qualitative analyses that serve as more specialized diagnostic tests. These in-depth analyses examine specific linguistic ‘symptoms’ to confirm the presence of TSD and understand its distinct manifestations (e.g., benign, insidious, overt, subtle). Just as an experienced doctor becomes progressively better at spotting diseases, TSD becomes easier to trace and critique once an analyst is sensitized to its discursive markers.
@@
For CDA to contribute to social change, however, it must engage in dialogue with social actors who are in a position to act. The ultimate value of this paper lies in fostering awareness among key stakeholders. This awareness, however, must be critical and extend beyond simply being able to trace TSD. It is also crucial for these social actors to understand why TSD may be problematic and how its components reinforce one another. There are three social groups that must develop this critical awareness for TSD to be challenged effectively: AI practitioners, policymakers, and consumers. AI practitioners building and deploying applications (e.g., engineers, designers) must be able to diagnose TSD in their discourse and that of their leadership. Without this critical awareness, even well-intentioned efforts can fall into a techno-solutionist trap. By becoming sensitized to TSD, AI practitioners can question whether a technological fix is the best approach and implement genuine solutions that account for complex social contexts and integrate appropriate safeguards. Policymakers and regulators formulating governance frameworks (e.g., the EU AI Act) are subject to intense lobbying and sophisticated public discourse. Being able to detect the subtle workings of TSD (e.g., distinguishing a genuine acknowledgment of risk from a token gesture meant to preempt regulation) is essential for crafting effective, robust, and truly society-serving policies that are not captured by the very industry they seek to govern. AI consumers (i.e., virtually everyone with access to digital technology) play a crucial, often underestimated role. If consumers are sensitized to TSD, they can purposefully exercise their consumer and civic power to question assumptions and claims that often dominate marketing and public narratives. Deciding whether to use AI and which one should be conscious and intentional, informed by the consumer’s own values.
@@
Developing a critical awareness to trace TSD is necessary but insufficient for meaningful praxis. While the diagnostic toolkit developed in this paper empowers social actors to identify and support more responsible manifestations (e.g., Benign Techno-Optimism) over more insidious or aggressive ones, the final stage of critical engagement requires holding elite members accountable by verifying the congruence between their public discourse and their material actions. Accountability demands that we scrutinize whether a company’s stated commitment to safety (e.g., Anthropic’s Responsible Scaling Policy) is reflected in its product design choices; whether its advocacy for responsible AI aligns with its lobbying efforts on regulation; and whether its vision of shared prosperity matches its internal labor practices and investment priorities. Without this vital step of verifying alignment between word and deed, even the most seemingly responsible discourse risks functioning merely as a strategic shield for unaccountable power.
@@ 6 Conclusion
This paper’s contributions should be contextualized by several limitations that offer potential avenues for future research. First, TSD may not be a consciously-held belief system. While this paper defines TSD as a cohesive ideology, its tenets might be more fragmented and/or opportunistic in practice. Future ethnographic or interview-based research could examine how well this framework aligns with the self-perceptions and motivations of elite members. Moreover, it could investigate the out-of-scope phenomena identified during the broad phase of this paper, including the role of saviorism given perceived moral motivations; internal disagreements and fragmentation within the elite (e.g., Andreessen versus Amodei on the merits of Universal Basic Income); and the quasi-religious nature of TSD discourse. A second limitation is the paper’s scope. The findings may not generalize to the broader global AI ecosystem, given the focus on a systematically defined, US-centric executive elite. Future analyses could include a broader range of influential actors, including leaders of non-US companies (e.g., DeepSeek, Mistral) and key opinion formers (e.g., scientists, journalists). These analyses would create a comparative understanding of how TSD is reproduced or challenged globally and across different institutional roles. Additionally, the scope could be broadened to include shorter texts (e.g., tweets) and spoken discourse (e.g., podcasts), thereby allowing the inclusion of highly influential figures such as Demis Hassabis and Elon Musk. Another limitation is that this paper does not establish a causal relationship between elite discourse and societal outcomes. While premised on the principle that discourse has social effects, it cannot isolate or measure the actual impact of these discursive strategies in the real world. This limitation points to the need for empirical reception studies. Future projects could employ experimental methods or surveys to investigate how different manifestations of TSD, from benign to insidious, actually influence public opinion and/or policymakers’ attitudes.
@@
The most personally exciting and immediate future project is the development of an AI-powered TSD diagnostic tool. The underlying model would be fine-tuned on this research’s annotations and findings and would be prompted to act as an expert in this paper’s ‘T’ methodology. The model could employ Retrieval-Augmented Generation (RAG) techniques to access TSD’s and CDA’s theoretical foundations and utilize Reinforcement Learning (RL) to incorporate user feedback on an ongoing basis. Such an application could analyze text, audio, and even video in near real-time, providing users with TSDA and TSDB scores, and identify concrete linguistic features that signal TSD’s presence. To maximize its practical impact, the tool could position analyzed texts on the TSDA/TSDB spectrum relative to this paper’s corpus and, crucially, link identified TSD aspects to established critiques and concrete examples of potential or actual harm. By presenting these insights in accessible formats (e.g., infographics, short videos, podcasts), this tool could translate critical academic research into a powerful instrument for public and professional awareness. This project would represent a reflexive application of AI, turning its analytical power back upon the ideological systems from which it emerges. This approach moves beyond a simple anti-technology stance to embody critical praxis. By leveraging AI’s unique capabilities for scalable, real-time analysis, the goal is not to propose a technological ‘fix’ for TSD, but to create an effective instrument for public sensitization.
@@
This paper is an anti-TSD piece of discourse from a concerned techno-optimist. While technology can be a powerful aid in tackling complex societal issues, it must be employed consciously to prevent or mitigate harm equitably. The sheer velocity of AI development with its accompanying legitimizing discourse makes critical awareness an immediate necessity. While awareness alone does not constitute transformative action, it is an essential prerequisite. By uncovering the Techno-Supremacy Doctrine’s core components, discussing its risks, and providing a diagnostic toolkit for its manifestations, this paper aims to foster a collective sensitization among AI practitioners, policymakers, and consumers. We all have the responsibility to act. Let us support or challenge dominant discourses both consciously and materially, demanding that the promises made are met with accountable action. Let us navigate the future of AI as active participants, not passive recipients of an “inevitable” destiny manufactured unilaterally by the AI executive elite.